An investor believes that the stock of the ABC corporation will remain in a trading range over the next 6 months, not trading above 70 or below 65 for this time period. ABC stock is now trading for 67 ½ and there are out of the money puts and calls available. The six month 65 put sells for 4 ¼ and the six month 70 call sells for 4. Describe a strategy for trading this opportunity and be specific about profitability and risk.
Answer #1

The stock of ABC Corporation trades between $ 70 and $ 65, with the current price being $ 67.5. A call with strike $ 70 sells for $ 4 and another put with strike $ 65 sells for $ 4,25. When a stock price is range bound in such a manner profit can be made by following the strategy given below:

- Sell the put option as well as the call option. The cash proceeds received will be (4 + 4.25) = $ 8.25

- The put position will be exercise by the put buyer only when the stock price goes below $ 65 which is unlikely to happen. Similarly, the call position will be exercised by the call buyer only when the stock price goes above $ 70 which again is unlikely to happen. This is so because the stock price is predicted to be range-bound between $ 65 and $ 70, thereby keeping both options out-of-the-money and making a clean profit for the option seller.

- The put position faces a risk of loss once, the stock price goes below the strike price of $ 65 and the call position faces a loss if and when the stock price goes above the strike price of $70. In both cases, an overall loss would be faced if the loss incurred is greater than the cash proceeds received by selling the option which is $ 8.25. Hence, put position faces a loss if stock price < (65 - 8.25) = $ 56.75 and the call position faces a loss if stock price > (70 + 8.25) = $ 78.25. For all stock prices between $56.75 and $ 78.25, the options seller makes a net profit.